There are 22 students in a music class in a school. The students in the class are either Freshmen or Sophomores, and there are 6 more boys than girls. If there are 5 Freshmen boys and 4 Sophomore girls, how many Sophomores are in the class?
Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(53763)

.

I just solved this problem at this forum a week ago. See the link

https://www.algebra.com/algebra/homework/word/misc/Miscellaneous_Word_Problems.faq.question.1201990.html


For your convenience, I place this my solution here again:


                                - - - - - - - - - - - -


It can be solved in three simple steps.


             Step 1 - determine how many boys are in the class


Let "b" be the number of boys in the class. 
Then the number of girls is (b-6).


The total students equation is

    b + (b-6) = 22,

which gives

    2b = 22 + 6,  2b = 28,  b = 28/2 = 14.  So, there are 14 boys in the class, in all.



             Step 2 - determine how many boys are the Sophomores


From total 14 boys, subtract 5 Freshmen boys and get 14-5 = 9 Sophomores boys.
   


             Step 3 - determine how many Sophomores are in the class


Add 4 Sophomore girls and 9 Sophomores boys to get 4+9 = 13 Sophomores in the class, in total.


ANSWER.  There are 13 Sophomores in the class, in all.

Solved.



Answer by Theo(13342)

i get 13 sophomores.

you are given that there are 22 students total.
you are given that there are 6 more boys than girls.
you are given that there are freshmen boys and 4 sophomore girls.
let b = number of boys and g = number of girls.
you get b + g = 22
since b = g + 6, you get g + 6 + g = 22.
combine like terms to get 2g + 6 = 22
subtract 6 from both sides of the equation to get 2g = 16
solve for g to get g = 8.
since b = g + 6, then b = 14
since 5 boys are freshmen, then 9 boys must be sophomores.
since 4 girls are sophomores, then 4 girls must be freshmen.
the freshman class has 5 boys and 4 girls = 9 total.
the sophomore class has 9 boys and 4 girls = 13 total.
the number of sophomores in the class is equal to 13.
i made a chart that you see below:
             freshmen    sophomores    total
boys             5           9           14
girls            4           4            8
total            9          13           22

looks good to me.
hopefully you agree.